# Sample Selection Bias in Machine Learning for Healthcare

### v1

@@ Abstract
While machine learning algorithms hold promise for personalised medicine, their clinical adoption remains limited. One critical factor contributing to this restraint is sample selection bias (SSB) which refers to the study population being less representative of the target population, leading to biased and potentially harmful decisions. Despite being well-known in the literature, SSB remains scarcely studied in machine learning for healthcare. Moreover, the existing techniques try to correct the bias by balancing distributions between the study and the target populations, which may result in a loss of predictive performance. To address these problems, our study illustrates the potential risks associated with SSB by examining SSB’s impact on the performance of machine learning algorithms. Most importantly, we propose a new research direction for addressing SSB, based on the target population identification rather than the bias correction. Specifically, we propose two independent networks (T-Net) and a multitasking network (MT-Net) for addressing SSB, where one network/task identifies the target subpopulation which is representative of the study population and the second makes predictions for the identified subpopulation. Our empirical results with synthetic and semi-synthetic datasets highlight that SSB can lead to a large drop in the performance of an algorithm for the target population as compared with the study population, as well as a substantial difference in the performance for the target subpopulations that are representative of the selected and the non-selected patients from the study population. Furthermore, our proposed techniques demonstrate robustness across various settings, including different dataset sizes, event rates, and selection rates, outperforming the existing bias correction techniques.
@@ 1 Introduction
Machine learning algorithms have demonstrated high diagnostic and prognostic accuracy in identifying and categorising diseases, promising personalised interventions and informed decision-making in healthcare [49, 58]. Their ability to analyse vast datasets and uncover hidden patterns surpasses traditional techniques and sometimes even human experts in task-specific applications, such as image-recognition tasks in radiology [36, 33]. However, despite their compelling potential and an increasing number of studies every year, their clinical adoption remains constrained by various issues [61]. One such issue is sample selection bias (SSB), recognised as a fundamental pitfall in the research design of clinical studies [65] but remains scarcely studied in machine learning for healthcare, posing a potential hurdle to its real-world applications.
Inherent in clinical studies, which serve as a source of data for machine learning, is the practice of sample selection, characterised by strict criteria for the inclusion and exclusion of patients [45]. However, these studies don’t add any bias as long as the study population is representative of the target population. The study population is defined as data used to develop the model and generally split into training, (internal) validation, and test datasets; and the target population is defined as the independent external data, i.e., deployment data or external validation data, respectively. SSB results from a non-uniform sampling process to select patients in the study population, leading to differences in the distributions of the study population and the target population [48, 6, 67, 7]. For example, in a clinical study, the outcomes of some patients are not available due to a loss to follow-up or competing events, and considering only those patients who completed the study introduces SSB in the dataset [2]. When machine learning algorithms are trained on such a biased dataset and employed on an unbiased dataset, it might lead to inaccurate predictions and suboptimal care for patient subpopulations not represented in the study population [50, 22]. Another example is the self-nomination of patients for a clinical study, which may not represent the entire population, such as the UK Biobank and might lead to SSB [56, 8]. Fig. 1 provides an example illustrating SSB and its impact on a prediction algorithm’s performance.
@@
SSB is well known for decades [25, 26, 3], and it has been discussed across several disciplines, including social science and econometrics [27, 28], environmental studies [62], finance [35], causality [31, 4], fairness [20], healthcare [44], and machine learning [66]. It is also an important consideration for clinical study design as the results of a biased study may not apply to a real patient population [6, 48]. Despite being well-known and a threat to the validity of algorithms [14], SSB has been scarcely studied in machine learning for healthcare [65]. This could be because (i) often less proportion of patients are excluded from the study population, making it hard to find enough data to correct the bias, (ii) most of the algorithms are not developed for the deployment, (iii) the bias is sometimes hidden under the assumed randomness in selection [51], (iv) risk estimation primarily belongs to causality, a principled approach to decision-making that examines cause-and-effect relationships. Causality has developed formal techniques, including causal diagrams, to address SSB [63, 53], and (v) SSB might not always lead to inaccurate predictions for the non-selected patients, as they might not influence the algorithm’s decision-making process.
It is not always possible to avoid SSB in the study but to utilise valuable information hidden in such a biased dataset, it is necessary to address the bias to ensure that decisions are made prudently and with reduced bias. Machine learning for healthcare lacked research to address SSB, however, the machine learning community has developed several techniques to address SSB [66] and to address distribution mismatch between the source and the target domains, such as covariate shift [52] or domain-adaptation [39], which have a setting similar to SSB. The existing techniques correct for the bias by aligning distributions of the study population and the target population, which could lead to a loss of predictive performance of the algorithm [66, 34, 39]. Moreover, in healthcare, the non-selected patients form a subpopulation that could greatly differ from the study population, making it hard to correct for the bias. Thus, the existing techniques may perform poorly for subpopulations that are representative of the non-selected subpopulation compared to those representative of the selected subpopulation.
@@
  • We propose a new research direction to address SSB in machine learning for healthcare, focusing on identification of the target subpopulation representative of the study population rather than the bias correction, and hence avoiding loss of predictive performance due to balancing of distributions.
@@ 3.1 SSB in Machine Learning for Healthcare
To address SSB in machine learning for healthcare, some studies have developed study-specific techniques. For example, Thakur et al., [57] identified selection bias in early-stage breast cancer prediction due to different observers scoring different regions of the same tumour. They proposed an automated method using whole-slide analysis to mitigate the bias. Mei and Xia, [43] considered selection bias due to the secondary use of regional electronic health records data and proposed a knowledge learning symbiosis framework to incorporate domain knowledge to address the bias. Additionally, in medical statistics and clinical studies, SSB is commonly treated as a missingness problem under missingness-at-random. Missing outcomes, along with other missing attributes of a patient, are imputed using multiple imputations. For example, Liu et al., [41] used imputation for post-menopausal breast cancer in the UK Biobank.
@@ 3.2 SSB in other Fields
From the above brief discussion, we find that SSB is well recognised as a threat to the generalisation of machine learning algorithms. Despite one of the major hurdles in the clinical adoption of machine learning algorithms, SSB is scarcely discussed. Moreover, the existing techniques to address SSB are based on bias correction which aligns the distributions between the study and the target populations and may lead to a loss of predictive performance of the prediction algorithms. Thus, there is a need to highlight the potential dangers of SSB in machine learning for healthcare and develop techniques that effectively leverage available data to provide accurate predictions, despite the presence of bias.
4 Target Population Identification
@@
In this section, we present the proposed target population identification (TPI) approach – a new research direction – to address SSB in machine learning for healthcare for developing algorithms that utilise the available biased data to make accurate predictions. The key idea of our proposed approach is that we constrain the underlying machine learning algorithms, at deployment time, to a subpopulation seen by the algorithm during the training process, i.e., to a target subpopulation which is representative of the study population, without requiring to balance distributions of the two populations. This differs from the existing SSB correction approaches which align the distributions of the study and the target populations, resulting in a loss of predictive performance of algorithms. Our novel approach is motivated by the observation that in healthcare the study and the target populations differ in terms of the non-selected patients. This is unlike domain adaptation where the target dataset is different to the source dataset and the aim is to perform better on the target dataset using information from the source dataset [39]. So for healthcare, an approach based on the identification of the target subpopulation representative of the study population can help in addressing SSB.
@@
Figure 2: An illustration of the proposed target population identification (TPI) approach which utilises all patients during training and employs a two-step evaluation process where first it identifies the target subpopulation representative of the study population, followed by predictions for the selected subpopulation.
@@
At training time (or development time), the proposed TPI approach, in addition to learning the underlying risk prediction task from the study population, learns the selection process of including patients in the study population using patient attributes, i.e., $P(s=1|\mathbf{x})$ using included as well as excluded patients. At test time (or deployment time), the proposed approach employs a two-step evaluation process. First, it identifies the target patient subpopulation that is representative of the study population, and then it applies the risk prediction algorithm to the patients identified in the first step, as depicted in Fig. 2. Thus, the proposed approach is based on the assumption that the selection process can be identified from patient characteristics $\mathbf{x}$. In other words, the proposed approach assumes that the outcomes of the non-selected patients are missing-at-random and it refers to case (ii) as discussed in Section 2. Moreover, the learning of the selection process has access to more data than the underlying risk prediction task because it considers included as well as excluded patients.
@@ 4.1 T-Net
To address SSB using the TPI approach, T-Net develops two independent neural networks, one for identification, i.e., to predict the selection/inclusion of a patient in the study population and the other network for the risk prediction task. The selection network trains a classifier $s=f:\mathbf{x}\rightarrow\{0,1\}$ over patients included as well as excluded from the study, where $s=1$ for patients included in the study population, i.e., $\mathbf{x}\in\mathbf{X}^{1}$ and $s=0$ for patients excluded from the study due to missing outcomes, i.e., $\mathbf{x}\in\mathbf{X}^{0}$. On the other hand, the risk prediction network trains a separate neural network $y=g:\mathbf{x}\rightarrow y$ over the study population (only patients included in the study), i.e., $\mathbf{x}\in\mathbf{X}^{1}$. T-Net employs a two-step evaluation where first the selection network identifies the target subpopulation belonging to the study population, followed by predictions by the risk predictor for the patients selected in the first step. Fig. 3(a) depicts the working of T-Net. Conversely, employing a naive machine learning algorithm would involve training the risk predictor $g(x)$ solely on data from selected patients, intending to make predictions for the entire target population. However, this approach may result in misclassifications for patients within the non-selected subpopulation, potentially leading to unnecessary or inadequate interventions. The risk predictor, developed on a biased dataset, lacks information about the excluded subpopulation, contributing to these shortcomings.
@@ 4.2 MT-Net
Therefore, both TPI approaches, T-Net and MT-Net, train on both the selected and the non-selected patients, employing a two-step evaluation. This ensures predictions are made for the target subpopulation encountered during training, yielding similar performances on both study and target populations—essential for the development and deployment of machine learning algorithms. Thus, the proposed approach not only offers a novel method for addressing SSB but also utilises the data which otherwise is not selected and is particularly crucial for data-limited studies. While T-Net provides flexibility with two separate networks, MT-Net benefits from shared learning, proving advantageous in settings with limited data for tackling SSB.
@@ 5.1 Experimental Setting
Baselines: We have utilised commonly used baselines, including IPW and imputation, and baselines from machine learning, including SSB and domain adaptation, as given below.
@@
Naive: As discussed in Section 4, a naive approach ignores SSB and develops an algorithm with a biased dataset and evaluates with an unbiased dataset as the algorithm does not have any mechanism to correct the bias or identify the target population. So, Naive should exhibit the lowest performance in the presence of SSB. Naive, Oracle and T-Net share an algorithm for risk prediction, however, the three differ in their evaluation process. Naive is evaluated on the entire test dataset, Oracle is evaluated only for patients belonging to the study population, and T-Net identifies the target subpopulation representing the study population and is evaluated only for the identified patients.
@@
Datasets: For synthetic data generation, we generated a dataset with 25 features, where features $\mathbf{X}$ are sampled from a uniform distribution over (-10, 10). For a uniformly sampled vector $\mathbf{a}$ of dimensions equal to features and a scalar $c$, outcomes $\mathbf{y}$ for risk prediction are generated using the following process,
@@
To add noise to outcomes, we randomly flip one percent of the outcomes. For introducing SSB to the dataset, we take another uniformly sampled vector $\mathbf{b}$ and a scalar $d$, and generate selection variable $\mathbf{s}$ using the following process,
@@
Hyperparameters and Performance Metrics: We employed neural networks for analysing SSB due to their flexibility, expressiveness, and popularity [13, 15]. We split the datasets as 20% of the full dataset as a test dataset, 20% of the remaining as a validation dataset and the rest as a training dataset. We utilised Adam optimiser [37] and tuned the learning rate from 0.0001 and 0.0005. A mini-batch size of 64 data points and 1000 epochs were used, and an early stopping with patience of 10 epochs. A dropout rate of 10% served as regularisation to prevent overfitting. The neural network architectures were tuned for hidden layers from [50], [100], [100, 100], where the number of list elements indicates the number of hidden layers, and values represent the number of neurons in the corresponding hidden layer. For task-specific layers, we tuned from [50], [100]. For a performance metric, we use the Area Under the Receiver Operating Characteristic Curve (AUC) to assess an algorithm’s ability to distinguish between positive and negative instances. It quantifies the area under the curve generated by plotting the true positive rate against the false positive rate at various classification thresholds. A higher AUC (closer to 1) indicates better discrimination performance, while a random classifier has an AUC of 0.5.

### v2

@@ Abstract
While machine learning algorithms hold promise for personalised medicine, their clinical adoption remains limited, partly due to biases that can compromise the reliability of predictions. In this paper, we focus on sample selection bias (SSB), a specific type of bias where the study population is less representative of the target population, leading to biased and potentially harmful decisions. Despite being well-known in the literature, SSB remains scarcely studied in machine learning for healthcare. Moreover, the existing machine learning techniques try to correct the bias mostly by balancing distributions between the study and the target populations, which may result in a loss of predictive performance. To address these problems, our study illustrates the potential risks associated with SSB by examining SSB’s impact on the performance of machine learning algorithms. Most importantly, we propose a new research direction for addressing SSB, based on the target population identification rather than the bias correction. Specifically, we propose two independent networks (T-Net) and a multitasking network (MT-Net) for addressing SSB, where one network/task identifies the target subpopulation which is representative of the study population and the second makes predictions for the identified subpopulation. Our empirical results with synthetic and semi-synthetic datasets highlight that SSB can lead to a large drop in the performance of an algorithm for the target population as compared with the study population, as well as a substantial difference in the performance for the target subpopulations that are representative of the selected and the non-selected patients from the study population. Furthermore, our proposed techniques demonstrate robustness across various settings, including different dataset sizes, event rates, and selection rates, outperforming the existing bias correction techniques.
@@ 1 Introduction
Machine learning algorithms have demonstrated high diagnostic and prognostic accuracy in identifying and categorising diseases, promising personalised interventions and informed decision-making in healthcare [53, 62]. Their ability to analyse vast datasets and uncover hidden patterns surpasses traditional techniques and sometimes even human experts in task-specific applications, such as image-recognition tasks in radiology [39, 36]. However, despite their compelling potential and an increasing number of studies every year, their clinical adoption remains constrained by various challenges [66], including biases that can undermine the validity of machine learning models. Among these, sample selection bias (SSB) has received limited attention in machine learning for healthcare, despite being recognised as a fundamental pitfall in the research design of clinical studies [70], posing a significant hurdle to real-world applications.
Inherent in clinical studies, which serve as a source of data for machine learning, is the practice of sample selection, characterised by strict criteria for the inclusion and exclusion of patients [49]. However, these studies don’t add any bias to results as long as the study population is representative of the target population. The study population is defined as data used to develop the model and generally split into training, (internal) validation and test datasets; and the target population is defined as the independent external data, i.e., deployment data or external validation data, respectively. SSB results from a non-uniform sampling process to select patients in the study population, leading to differences in the distributions of the study population and the target population [52, 6, 72, 7]. For example, suppose in a clinical study, the outcomes of some patients are not available due to a loss to follow-up or competing events, and considering only those patients who completed the study introduces SSB in the dataset [2]. When machine learning algorithms are trained on such a biased dataset and employed on an unbiased dataset, it might lead to inaccurate predictions and suboptimal care for patient subpopulations not represented in the study population [54, 23]. Another example is the self-nomination of patients for a clinical study, which may not represent the entire population, such as the UK Biobank and might lead to SSB [60, 8]. Fig. 1 provides an example illustrating SSB and its impact on a prediction algorithm’s performance.
@@
SSB is well known for decades [27, 29, 3], and it has been discussed across several disciplines, including social science and econometrics [30, 31], environmental studies [67], finance [38], causality [34, 4], fairness [21], healthcare [48], and machine learning [71]. It is also an important consideration for clinical study design as the results of a biased study may not apply to a real patient population [6, 52]. Despite being well-known and a threat to the validity of algorithms [15], SSB has been scarcely studied in machine learning for healthcare [70]. This could be potentially because (i) often less proportion of patients are excluded from the study population, making it hard to find enough data to correct the bias, (ii) most of the algorithms are not developed for the real-world deployment, (iii) the bias is sometimes hidden under the assumed randomness in selection [55], (iv) risk estimation primarily belongs to causality, a principled approach to decision-making that examines cause-and-effect relationships. Causality has developed formal techniques, including causal diagrams, to address SSB [68, 57], and (v) SSB might not always lead to inaccurate predictions for the non-selected patients, as they might not influence the algorithm’s decision-making process.
It is not always possible to avoid SSB in the study but to utilise valuable information hidden in such a biased dataset, it is necessary to address the bias to ensure that decisions are made prudently and with reduced bias. Machine learning for healthcare lacked research to address SSB, however, the machine learning community has developed several techniques to address SSB [71] and to address distribution mismatch between the source and the target domains, such as covariate shift [56] or domain-adaptation [42], which have a setting similar to SSB. The existing machine learning techniques correct for the bias by aligning distributions of the study population and the target population, which could lead to a loss of predictive performance of the algorithm [71, 37, 42]. Moreover, in healthcare, the non-selected patients form a subpopulation that could greatly differ from the study population, making it hard to correct for the bias. Thus, the existing techniques may perform poorly for subpopulations that are representative of the non-selected subpopulation compared to those representative of the selected subpopulation.
@@
  • We propose a new research direction to address SSB in machine learning for healthcare, focusing on identification of the target subpopulation representative of the study population rather than the bias correction, and hence avoiding loss of predictive performance due to aligning distributions between the study and the target populations.
@@ 3.1 SSB in Machine Learning for Healthcare
To address SSB in machine learning for healthcare, some studies have developed study-specific techniques. For example, Thakur et al., [61] identified selection bias in early-stage breast cancer prediction due to different observers scoring different regions of the same tumour. They proposed an automated method using whole-slide analysis to mitigate the bias. Similarly, Mei and Xia, [46] considered selection bias due to the secondary use of regional electronic health records data and proposed a knowledge learning symbiosis framework to incorporate domain knowledge to address the bias. Additionally, in medical statistics and clinical studies, SSB is commonly treated as a missingness problem under missingness-at-random. Missing outcomes, along with other missing attributes of a patient, are imputed using multiple imputations. For example, Liu et al., [44] used imputation for post-menopausal breast cancer in the UK Biobank.
@@ 3.2 SSB in other Fields
From the above brief discussion, we find that SSB is well recognised as a threat to the generalisation of machine learning algorithms. Despite one of the major hurdles in the clinical adoption of machine learning algorithms, SSB is scarcely discussed. Moreover, the existing techniques to address SSB are based on bias correction which forces alignment of the distributions between the study and the target populations and may lead to a loss of predictive performance of the prediction algorithms. Thus, there is a need to highlight the potential dangers of SSB in machine learning for healthcare to encourage further research and develop techniques that effectively leverage available data to provide accurate predictions, despite the presence of bias.
4 Methods
@@
In this section, we present the proposed target population identification (TPI) approach – a new research direction – to address SSB in machine learning for healthcare for developing algorithms that utilise the available biased data to make accurate predictions. The key idea of our proposed approach is that we constrain the underlying machine learning algorithms, at deployment time, to a subpopulation seen by the algorithm during the training process, i.e., to a target subpopulation which is representative of the study population, without requiring to balance distributions of the two populations. This differs from the existing SSB correction approaches which forces alignment of the distributions of the study and the target populations, resulting in a loss of predictive performance of algorithms. Our novel approach is motivated by the observation that in healthcare the study and the target populations differ in terms of the non-selected patients. This is unlike domain adaptation where the target dataset is different to the source dataset and the aim is to perform better on the target dataset using information from the source dataset [42]. So for healthcare, an approach based on the identification of the target subpopulation representative of the study population can help in addressing SSB.
@@
Figure 2: An illustration of the proposed target population identification (TPI) approach which employs a two-step evaluation process where first it identifies the target subpopulation representative of the study population, followed by predictions for the selected subpopulation.
@@
At training time (or development time), the proposed TPI approach, in addition to learning the underlying risk prediction task from the study population, also learns the selection process of including patients in the study population using patient attributes, i.e., $P(s=1|\mathbf{x})$ using included as well as excluded patients. At test time (or deployment time), the proposed approach employs a two-step evaluation process. First, it identifies the target patient subpopulation that is representative of the study population, and then it applies the risk prediction algorithm to the patients identified in the first step, as depicted in Fig. 2. Thus, the proposed approach is based on the assumption that the selection process can be identified from patient characteristics $\mathbf{x}$. In other words, the proposed approach assumes that the outcomes of the non-selected patients are missing-at-random and it refers to case (ii) as discussed in Section 2. Moreover, the learning of the selection process has access to more data than the underlying risk prediction task because it considers included as well as excluded patients.
@@ 4.1 T-Net
To address SSB using the TPI approach, T-Net develops two independent neural networks, one for identification, i.e., to predict the selection/inclusion of a patient in the study population and the other network for the risk prediction task. The two networks can be trained in any order. The selection network trains a classifier $s=f:\mathbf{x}\rightarrow\{0,1\}$ over patients included as well as excluded from the study, where $s=1$ for patients included in the study population, i.e., $\mathbf{x}\in\mathbf{X}^{1}$ and $s=0$ for patients excluded from the study due to missing outcomes, i.e., $\mathbf{x}\in\mathbf{X}^{0}$. On the other hand, the risk prediction network trains a separate neural network $y=g:\mathbf{x}\rightarrow y$ over the study population (only patients included in the study), i.e., $\mathbf{x}\in\mathbf{X}^{1}$. T-Net employs a two-step evaluation where first the selection network identifies the target subpopulation belonging to the study population, followed by predictions by the risk predictor for the patients selected in the first step. Fig. 3(a) depicts the working of T-Net. Conversely, employing a naive machine learning algorithm would involve training the risk predictor $g(x)$ solely on data from selected patients, intending to make predictions for the entire target population. However, this approach may result in misclassifications for patients within the non-selected subpopulation, potentially leading to unnecessary or inadequate interventions. The risk predictor, developed on a biased dataset, lacks information about the excluded subpopulation, contributing to these shortcomings.
@@ 4.2 MT-Net
Therefore, both TPI approaches, T-Net and MT-Net, train on both the selected and the non-selected patients, and employ a two-step evaluation. This ensures predictions are made for the target subpopulation encountered during training, yielding similar performances on both study and target populations – essential for the development and deployment of machine learning algorithms. Thus, the proposed approach not only offers a novel method for addressing SSB but also utilises the data which otherwise is not selected and is particularly crucial for data-limited studies. While T-Net provides flexibility with two separate networks, MT-Net benefits from shared learning, proving advantages in settings with limited data for tackling SSB. The architectures of T-Net and MT-Net follow standard single-task and multitasking networks in deep learning, respectively. The number of layers and neurons in each of the layers are described in Section 5.1 under the heading ‘Hyperparameters and Performance Metrics’.
@@ 5.1 Experimental Setting
Baselines: We have utilised commonly used baselines, including IPW and imputation, and baselines from machine learning, including SSB handling and domain adaptation techniques, as given below.
@@
Naive: As discussed in Section 4, a naive approach ignores SSB and develops an algorithm with a biased dataset and evaluates with an unbiased dataset as the algorithm does not have any mechanism to correct the bias or identify the target population. So, Naive should exhibit the lowest performance in the presence of SSB. Naive, Oracle and T-Net methods have same shared neural network for risk prediction, however, the three differ in their evaluation process. Naive is evaluated on the entire test dataset, Oracle is evaluated only for patients belonging to the study population, and T-Net identifies the target subpopulation representing the study population and is evaluated only for the identified patients.
@@
Datasets: For synthetic data generation, following the literature [11], we generated a dataset with 25 features, slightly more than in real-world datasets discussed below, where features $\mathbf{X}$ are sampled from a uniform distribution over (-10, 10). For a uniformly sampled vector $\mathbf{a}$ of dimensions equal to features and a scalar $c$, outcomes $\mathbf{y}$ for risk prediction are generated using the following process,
@@
To add noise to outcomes, we randomly flip one percent of the outcomes to reflect real-world data quality issues and imperfections. For introducing SSB to the dataset, we take another uniformly sampled vector $\mathbf{b}$ and a scalar $d$, and generate selection variable $\mathbf{s}$ using the following process,
@@
Hyperparameters and Performance Metrics: We employed neural networks for analysing SSB due to their flexibility, expressiveness, and popularity [14, 16], and the fact that SSB handling methods are mostly implemented with neural networks [37, 24] and there is no direct way to implement multitasking with standard machine learning algorithms. We split the datasets as 20% of the full dataset as a test dataset, 20% of the remaining as a validation dataset and the rest as a training dataset. We utilised Adam optimiser [40] and tuned the learning rate from 0.0001 and 0.0005. A mini-batch size of 64 data points and 1000 epochs were used, and an early stopping with patience of 10 epochs. A dropout rate of 10% served as regularisation to prevent overfitting. For all the methods, the neural network architectures were tuned for hidden layers from [50], [100], [100, 100], where the number of list elements indicates the number of hidden layers, and values represent the number of neurons in the corresponding hidden layer. For task-specific layers, we tuned from [50], [100]. For a performance metric, we use the Area Under the Receiver Operating Characteristic Curve (AUC) to assess an algorithm’s ability to distinguish between positive and negative instances. It quantifies the area under the curve generated by plotting the true positive rate against the false positive rate at various classification thresholds. A higher AUC (closer to 1) indicates better discrimination performance, while a random classifier has an AUC of 0.5.